### v1

# Magnetic-field-tuned randomness in inhomogeneous altermagnets

Abstract
Altermagnetic (AM) states have compensated collinear magnetic configurations that are invariant under a combination of real-space rotation and time reversal. While these symmetries forbid a direct bilinear coupling of the AM order parameter with a magnetic field, they generally enable piezomagnetism, manifested as a trilinear coupling with magnetic field and strain. Here, we show that, because of this coupling, in an altermagnet subjected to random strain, the magnetic field triggers an effective random field conjugate to the AM order parameter, providing a rare realization of a tunable random-field Ising model. Specifically, we find two competing effects promoted by an external magnetic field: an increasing random-field disorder, which suppresses long-range AM order, and an enhanced coupling to elastic fluctuations, which favors AM order. By solving the corresponding random-field transverse-field Ising model via a mean-field approach, we obtain the temperature-magnetic field phase diagram of an inhomogeneous AM state for different strengths of random-strain disorder, unveiling the emergence of a field-induced reentrant AM phase. We also discuss the fingerprints of this rich behavior on several experimentally-accessible quantities, such as the shear modulus, the elasto-caloric effect coefficient, and the AM order parameter. Our results reveal an unusual but experimentally-feasible path to tune AM order with uniform magnetic fields.
I Introduction
A recent classification of collinear magnetism revealed the existence of three different types of magnetic states: ferromagnetism, antiferromagnetism, and altermagnetism [1, 2, 3]. These different magnetic ground states, which are formally defined in the absence of spin-orbit coupling using the concept of spin groups [3, 4, 5, 6, 7], are distinguished according to the type of space group operation which, when combined with time-reversal, leaves the ground state invariant. In ferromagnets (FM), there is no such operation and a uniform splitting between spin-up and spin-down bands emerges. Time-reversed antiferromagnetic (AFM) ground states, on the other hand, are related by a translation or inversion, resulting in a symmetry-protected Kramers spin degeneracy throughout the entire Brillouin zone. Finally, time-reversed altermagnetic (AM) ground states are related by any crystalline operation that is not translation or inversion, such as rotation, mirror reflection, or non-symmorphic operations like glides or screw rotations. The implication of this symmetry is a nodal $d$-wave, $g$-wave, or $i$-wave spin-split band structure, with spin degeneracy preserved only along certain high-symmetry momentum space planes (for a recent review, see [8]). These unusual properties have motivated extensive theoretical works on the interplay between altermagnetism and other electronic phenomena such as topology [9, 10, 11, 12, 13, 14, 15, 16], electronic correlations [17, 18, 19, 20, 21, 22], superconductivity [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], non-trivial responses [38, 39, 40, 41, 42, 43, 44, 45, 46], and multiferroics [47, 48, 49].
[…]
Figure 1: Schematic phase diagram based on the results of this paper for an Ising-like altermagnet in the presence of weak (top) or strong (bottom) random strain disorder-strength as a function of magnetic field $H_{z}$, quantum fluctuations tuning parameter $\Gamma$, and temperature $T$. Above a critical disorder strength, the altermagnetic phase stabilizes into separate low-field and high-field ordered phases.
At first sight, one may think that a magnetic field $\mathbf{H}$ is not an ideal tuning parameter for altermagnetism. Indeed, since $\mathbf{H}$ cannot couple directly (i.e., via a bilinear coupling) to the AM order parameter $\boldsymbol{\Phi}(\mathbf{x})$, one would expect $\mathbf{H}$ to have a minimal effect on the onset of AM order. However, the symmetries that define altermagnetism also imply that nearly all altermagnets display piezomagnetism [39, 10, 42, 94, 95]. Piezomagnetism is a phenomenon analogous to piezoelectricity, in which the application of strain leads to a magnetic dipole moment, and vice-versa [96]. It is characterized by the linear response equation
$H_{i}=\Lambda_{ijk}\varepsilon_{jk}$ (1)
where $\Lambda_{ijk}$ is the piezomagnetic tensor, and $\varepsilon_{jk}$ is the strain tensor defined in terms of the crystal displacement field $\mathbf{u}$ through $\varepsilon_{jk}=\frac{1}{2}(\partial_{j}u_{k}+\partial_{k}u_{j})$. As discussed elsewhere, in altermagnets the components of $\Lambda_{ijk}$ are proportional to odd powers of the AM order parameter $\boldsymbol{\Phi}$ [46]. Moreover, $\varepsilon_{jk}$ is generally a shear strain, since a pure symmetry-preserving dilatation $\sum_{i}\varepsilon_{ii}$ does not couple to a product of magnetic field and an AM order parameter. We emphasize that piezomagnetism is a qualitatively different response than magnetostriction, which relates strain $\varepsilon_{ij}$ to a magnetic field bilinear $H_{j}H_{k}$ [97, 98]. Because of piezomagnetism, the Landau free energy of an altermagnet must have a term of the form
$\mathcal{F}_{\text{pzm}}=-\lambda_{ijkl}\Phi_{i}H_{j}\varepsilon_{kl}$ (2)
where the coupling constants $\lambda_{ijkl}$ are related to the piezomagnetic tensor. In this paper, we show that, because of piezomagnetism, a magnetic field can be used to tune the AM transition, provided that the system displays an inhomogeneous distribution of internal strain fields. This is generally expected to be the case in any crystal, since unavoidable lattice defects such as dislocations, vacancies, and dopants always generate random strain $\varepsilon_{ij}(\mathbf{x})$. The key point is that, from Eq. (2), the product of the appropriate component of $\mathbf{H}$ with an inhomogeneous strain field $\varepsilon_{ij}(\mathbf{x})$ acts as an effective random longitudinal field that is conjugate to $\boldsymbol{\Phi}$. Thus, in the usual case in which spin-orbit coupling (SOC) lowers the symmetry of the vector AM order parameter to a single component $\Phi$ [10], the inhomogeneous AM realizes a rare example of a tunable random field Ising model (RFIM). The RFIM is the prototypical model to elucidate the impact of disorder on critical phenomena [99, 100, 101, 102, 103]. The tuning parameter of the model is the disorder strength, set by, e.g., the width of the longitudinal field distribution. While the RFIM is realized in certain magnetic [104, 105] and nematic materials [106], the disorder strength is usually (but not always [107]) fixed for a given crystal. In contrast, in the case of this RFIM realization in altermagnets, the disorder strength is continuously tuned by the magnetic field even though the distribution of inhomogeneous strain is unchanged.
We show that the random-field effect generated by the combination of magnetic fields and residual strain, which tends to suppress AM order, competes with another effect that also arises from piezomagnetism but that tends to favor AM order. This latter effect arises because, in the presence of an external field, thermally excited elastic fluctuations mediate long-range correlations between the AM degrees of freedom. To investigate the interplay between these two piezomagnetic-generated effects, we employ a mean-field approach to calculate the temperature-magnetic field phase diagram of an inhomogeneous AM state. We find two qualitatively distinct behaviors, illustrated in Fig. 1. For weak disorder strength, application of a magnetic field first suppresses the AM transition temperature but then enhances it. For large disorder strength, the random-field effect of the magnetic field is strong enough to completely suppress the AM phase for intermediate field values, leading to a guaranteed reentrance behavior for large enough fields. These behaviors are also reflected at $T=0$, where a non-thermal tuning parameter (denoted by $\Gamma$ in the phase diagrams) tunes a quantum AM transition. We also compute the behavior of several experimentally observable quantities across these phase diagrams, such as the AM order parameter, the shear modulus, and the elasto-caloric effect coefficient, thus providing concrete experimental predictions to verify this behavior in candidate AM materials.
This paper is organized as follows. In Sec. II we construct an effective random-field transverse-field Ising model (RF-TFIM) for an AM system [108]. This model consists of a uniform transverse field $\Gamma$ encapsulating the role of quantum fluctuations, and a longitudinal field $H_{z}$ that promotes the coupling between $\Phi$ and shear strain $\varepsilon$. We compute the mean-field phase diagram of this model as a function of the parameters $T,$ $\Gamma$, $H_{z}$, and intrinsic random strain. In Sec. III we compute the order parameter self-consistently and obtain the AM susceptibility, shear modulus renormalization, and elasto-caloric effect both as a function of $T$ and of $H_{z}$. Like in the case a nematic critical point, we find a softening of the shear modulus at the AM critical point signaling an accompanying structural transition. In Sec. IV, we discuss the effect of fluctuations beyond mean-field, which are typical for random field models without infinite-range interactions. Finally, we present our conclusions in Sec. V, and in the Appendix, we provide additional details of the calculation for the elasto-caloric effect coefficient.
II Random-field Ising model with a uniform transverse field
[…]
$\mathcal{F}_{\text{pzm}}=-\lambda\Phi H_{z}\varepsilon$ (3)
where $\lambda$ is the piezomagnetic coupling, $\varepsilon\equiv\varepsilon_{xy}$ in the case of a $d_{xy}$-wave AM and $\varepsilon\equiv\varepsilon_{x^{2}-y^{2}}$ in the case of a $d_{x^{2}-y^{2}}$-wave AM.
To proceed, we write down an effective low-energy model for the coupled AM-strain degrees of freedom. Since the AM order parameter is Ising like, we model it in terms of a transverse-field Ising model Hamiltonian:
$\mathcal{H}_{\mathrm{AM}}=-J\sum_{\langle ij\rangle}\tau_{i}^{z}\tau_{j}^{z}-\Gamma\sum_{i}\tau_{i}^{x}$ (4)
Here, $\tau_{i}^{z}$ and $\tau_{i}^{x}$ are Pauli matrices, where the thermal average $\langle\tau_{i}^{z}\rangle$ is the AM order parameter. $J$ is an effective AM interaction that sets the scale of the thermal AM transition (and should not be confused with an exchange interaction) whereas the transverse field $\Gamma$ is a non-thermal parameter (such as doping or pressure) that promotes quantum fluctuations, i.e., tunneling between local pseudo-spin states. Thus, as $\Gamma$ is enhanced, the ordered state is suppressed and the system is driven to a quantum critical point (QCP).
Strain fields in crystals are usually inhomogeneous, $\varepsilon\to\varepsilon(\mathbf{x})$, with two contributions arising from distinct phenomena. The first is a non-singular contribution from thermally excited elastic fluctuations $\varepsilon_{0}(\mathbf{q})$, associated with the shear modulus $C_{0}$ defined as $C_{0}\equiv C_{66}$ for a $d_{xy}$-wave AM and $C_{0}\equiv(C_{11}-C_{12})/2$ for a $d_{x^{2}-y^{2}}$-wave AM. The second is a singular contribution from random strain arising from crystal defects $\varepsilon_{s}(\mathbf{x})$ in the form of a quenched random field. Therefore, from Eq. (3), we can cast the inhomogeneous piezomagnetic term as:
$\mathcal{H}_{\text{pzm}}=-\lambda H_{z}\sum_{i}(\varepsilon_{0,i}+\varepsilon_{s,i})\tau_{i}^{z}$ (5)
Note that the elastic fluctuations are well-defined excitations; therefore $\varepsilon_{0,i}$ should be thought of as an annealed dynamical field which one can integrate out. The effect of the non-homogeneous part of $\varepsilon_{0,i}$, i.e., the $\mathbf{q}\neq 0$ modes, is to drive the AM transition mean field and renormalize the effective Hamiltonian [110, 111]. We therefore dispense with $\varepsilon_{0}(\mathbf{q}\neq 0$) modes and regard them as subleading corrections to the full Hamiltonian. By contrast, $\varepsilon_{s,i}$ represents quenched disorder and remains as a fixed realization of a random field drawn from a zero-mean probability distribution. Putting it all together, the effective Hamiltonian consisting of altermagnetic and elastic degrees of freedom is
$\mathcal{H}$ $=-J\sum_{\langle ij\rangle}\tau_{i}^{z}\tau_{j}^{z}-\frac{N}{2}C_{0}\varepsilon_{0}^{2}-\lambda H_{z}\sum_{i}(\varepsilon_{0}+\varepsilon_{s,i})\tau_{i}^{z}$ (6)
$-\Gamma\sum_{i}\tau_{i}^{x}$
where $N$ denotes the number of sites.
III Field-tuned altermagnetic transition
III.1 Renormalized field-dependent Hamiltonian
In this section, we analyze the phase diagram of Eq. (6) within mean-field over a wide range of parameters, including the strength of quantum fluctuations, temperature, disorder strength, and magnetic field. To formally perform a mean-field calculation, we extend the AM interaction $J$ to also be infinite-range rather than nearest-neighbors:
$-J\sum_{\langle ij\rangle}\to-\frac{J}{N}\sum_{i<j}$ (7)
Next, we integrate out the uniform elastic field $\varepsilon_{0}$, which introduces an infinite-range interaction, effectively enhancing $J\to\tilde{J}$ for nonzero $H_{z}$. This yields a renormalized Hamiltonian of the form
~$\tilde{\mathcal{H}}=-\frac{\tilde{J}}{N}\sum_{i<j}\tau_{i}^{z}\tau_{j}^{z}-\lambda H_{z}\sum_{i}\varepsilon_{s,i}\tau_{i}^{z}-\Gamma\sum_{i}\tau_{i}^{x}$ (8)
where
$\tilde{J}\equiv J[1+(H_{z}/H_{\lambda})^{2}]$ (9)
Here, we introduced a magnetic field scale $H_{\lambda}$ defined as
$H_{\lambda}\equiv\frac{\sqrt{C_{0}J}}{\lambda}$ (10)
These results follow from a simple Gaussian identity applied to the $\varepsilon_{0}$-dependent part of $\mathcal{H}$:
$\begin{split}-T\log\int&\mathrm{d}\varepsilon_{0}e^{-\frac{N}{2T}C_{0}\varepsilon_{0}^{2}+\frac{\lambda H_{z}}{T}\varepsilon_{0}\sum_{i}\tau_{i}^{z}}\\ &=-\frac{1}{N}\Big{(}\frac{H_{z}}{H_{\lambda}}\Big{)}^{2}\sum_{i<j}\tau_{i}^{z}\tau_{j}^{z}\end{split}$ (11)
and further justify replacing the original nearest-neighbor interaction with an all-to-all interaction.
[…]
We now make a few convenient substitutions. First, we express all energy scales in units of $J$ by setting $J=1.$ Second, we express the Hamiltonian in a dimensionless form by dividing through by the $H_{z}$-dependent renormalized exchange $\tilde{J}=1+(H_{z}/H_{\lambda})^{2}$ and arrive at
~$\tilde{\mathcal{H}}/\tilde{J}=-\frac{1}{N}\sum_{i<j}\tau_{i}^{z}\tau_{j}^{z}-\sum_{i}z_{i}\tau_{i}^{z}-\gamma\sum_{i}\tau_{i}^{x}$ (12)
Here, $z_{i}$ and $\gamma$ are effective longitudinal (random) and transverse (uniform) fields, which take into account the exchange enhancement due to elastic fluctuations. By construction, they are $H_{z}$-dependent. Upon defining a dimensionless magnetic field parameter $h=H_{z}/H_{\lambda}$, the quantity $z_{i}$ takes the form
$z_{i}=z_{i}(h)$ $=\frac{2h}{1+h^{2}}\left(\frac{\lambda H_{\lambda}\varepsilon_{s,i}}{2}\right)$ (13)
It is clear that the random field $z_{i}(h)$ reaches its maximum value at $h=1$, i.e., at $H_{z}=H_{\lambda}$, when it is equal to the term inside the brackets. For concreteness, we regard $\varepsilon_{s,i}$ as Gaussian-distributed strain, with standard deviation $\bar{\varepsilon}_{s}$ such that $z_{i}$ also follows a Gaussian distribution with $h$-dependent standard deviation
$w(h)=\frac{2h}{1+h^{2}}W$ (14)
where
$W$ $=\frac{\lambda H_{\lambda}\bar{\varepsilon}_{s}}{2}\equiv\sqrt{\frac{E_{\text{el}}}{2}}$ (15)
Here
$E_{\text{el}}=\frac{C_{0}\bar{\varepsilon}_{s}^{2}}{2}$ (16)
is a dimensionless quantity that gives the elastic energy of the random distribution of strain in units of the AM interaction $J$. In what follows, the standard deviation of the random variables $W$ shall be regarded as a parameter of the problem. Similarly, the temperature and transverse field in these dimensionless units are
$t(h)$ $=\frac{T}{1+h^{2}}$ (17)
$\gamma(h)$ $=\frac{\Gamma}{1+h^{2}}$ (18)
where, we recall, $T$ and $\Gamma$ are given in units of $J$.
To set the stage, we provide a qualitative description of the phase diagram in the four-dimensional $(h,W,\Gamma,T)$ parameter space. In the classical limit $\Gamma=0$, Eq. (12) describes the well-known infinite-range RFIM, for which there exists a critical random field strength $w_{c}$ separating an ordered and a disordered phase [100]. Unlike the prototypical RFIM, however, here the random field strength $w(h)$ is tunable by magnetic field, suggesting the possibility of field values $h_{1}$ and $h_{2}$ for which $w(h_{1})<w_{c}$ and $w(h_{2})>w_{c}$. This implies the existence of a critical field $h_{c}\in[h_{1},h_{2}]$ marked by $w(h_{c})=w_{c}$ which separates the ordered and disordered phases. Additionally, $w(h)$ is non-monotonic, and reaches a maximum value of $W\propto\bar{\varepsilon}_{s}$ at $h=1$. Therefore three possibilities exist when the system begins in the ordered phase and a magnetic field is applied, which we delineate as follows:
  • In the regime of strong disorder, $W$ exceeds $w_{c}$. The non-monotonicity of $w(h)$ implies the existence of two critical fields $h_{c}^{-}<1$ and $h_{c}^{+}>1$, both satisfying $w(h_{c}^{\pm})=w_{c}$, for which the system is disordered over a finite window $[h_{c}^{-},h_{c}^{+}]$. This implies the phenomenon of altermagnetic reentrance, in which the AM order parameter $\Phi$ is nonzero for $h<h_{c}^{-}$ and $h>h_{c}^{+}$ and is identically zero for $h\in[h_{c}^{-},h_{c}^{+}]$.
  • In the regime of weak disorder, $W$ is smaller than $w_{c}$. As a result, the system remains trapped in the ordered phase. Interestingly, the order parameter $\Phi$ exhibits a vestige of the reentrance behavior characteristic of the strong disorder case, as it attains a non-zero minimum value $\Phi_{\text{min}}$ at a field value $h_{\text{min}}<1$ for arbitrarily small $W\propto\bar{\varepsilon}_{s}$.
  • In the marginal disorder regime, $W=w_{c}$. As a result, $w(h)=w_{c}$ exhibits a double root $h=h_{*}$. Given that $h_{*}$ is a double root, tuning the magnetic field across $h_{*}$ amounts to running tangent to the co-dimension $1$ phase boundary at precisely one point ($h_{*},W_{*},\Gamma_{*},T_{*}$). The process of taking $W$ from $W<w_{c}$ to $W=w_{c}$ causes $\Phi(h)$ to smoothly deform with $\Phi_{\text{min}}$ going to zero at a single point $h_{*}$. This gives rise to a “V”-like behavior of the order parameter, $\Phi\propto|h-h_{*}|$. This unusual phenomenon, which we explore in Section III, reflects the nontrivial RF-TFIM parameter space enabled by the piezomagnetic coupling.
The presence of thermal and quantum fluctuations reduces the critical strength $w_{c}$, i.e., increases the effective disorder. Formally, this means that the function $w_{c}(\Gamma,T)$ decreases with increasing $\Gamma$ and $T$, and consequently, the reentrant behavior is enhanced, as the paramagnetic window $\Delta h=h_{c}^{+}-h_{c}^{-}$ grows. Because $W$ is an intrinsic property of the crystal, it is unlikely to be tunable. Therefore, to assess the features of the phase diagram delineated above one must tune $\Gamma$ and $T$ so that the function $w_{c}(\Gamma,T)$ becomes greater than, equal to, or less than the intrinsic random strain scale $W$.
III.2 Zero-temperature phase diagram
Figure 2: Zero-temperature mean-field phase diagram of the inhomogeneous altermagnet as a function of relative disorder strength $W=\sqrt{E_{\text{el}}/2}$, quantum-fluctuations promoting transverse-field $\Gamma$, and scaled magnetic field $h\equiv H_{z}/H_{\lambda}$. For $W>W_{*}(\Gamma)$, a horizontal line intersects the AM-PM critical surface twice, corresponding to AM reentrance. We show constant-$W$ cuts of this plot in Fig. 3.
Figure 3: (a): $T=0$ phase diagram with magnetic-field scaled axes $\gamma\equiv\Gamma/(1+h^{2})$ and $w\equiv 2Wh/(1+h^{2})$, where $\Gamma=0.8J$ sets the starting point of the phase trajectory along the $\gamma$-axis and $W$ (listed in the legend) sets the maximum extent of the trajectories along the $w$-axis. Increasing the magnetic field moves the system along the trajectory as indicated by the arrows. (b-e): $T=0$ phase diagram with unscaled axes $h=H_{z}/H_{\lambda}$ and $\Gamma$ corresponding to $W/W_{0}=0.4$, $0.9$, $1.0$, and $1.05$, respectively. As the disorder strength $W$ increases, the paramagnetic region grows and the AM-PM transition line forms a bulge at $(\Gamma_{*},h_{*})$ that annihilates at $(0,1)$.
We now construct the mean-field $T=0$ phase diagram of Eq. (8) as a function of $h$, $\Gamma$, and $W$, which is shown in Fig. 2. In preparation for the finite temperature case, for which there are four parameters ($h,W,\Gamma,T$) rather than three, we describe two complementary approaches to visualize the phase diagram. The first is to construct a single 2D phase diagram with $h$-dependent axes $w(h)$ and $\gamma(h)$, where the AM-PM phase boundary appears as a convex curve bounding the origin (Fig. 3a). Tuning $h$ amounts to following an $h$-parameterized curved trajectory in this ($w,\gamma$)-plane, which may or may not intersect with the phase boundary depending on the values of $W$ and $\Gamma$. The second approach is to construct multiple 2D phase diagrams, one for each $W$ – corresponding to different strengths of random strain – with $h$ and $\Gamma$ as dimensionless axes (Fig. 3b-e). In this case, tuning away from the $h=0$ limit amounts to following a line parallel to the $h$-axis. The appearance of reentrance over some parameter regime manifests as a “bulge” in the AM-PM phase boundary. Finally, a third, which is relevant only for the $T=0$ case, is to represent the full ($h,W,\Gamma$) phase diagram, as depicted in Fig. 2.
We delineate the first approach now. The $h$-parameterized trajectory $\boldsymbol{\theta}$ in this ($w,\gamma$)-plane is given by
$\boldsymbol{\theta}(W,\Gamma;h)=\langle w(h),\gamma(h)\rangle=\frac{\langle 2hW,\Gamma\rangle}{1+h^{2}}$ (19)
Note the special values
$\begin{cases}\boldsymbol{\theta}(0)&=\langle 0,\Gamma\rangle\\ \boldsymbol{\theta}(1)&=\langle W,\Gamma/2\rangle\\ \boldsymbol{\theta}(\infty)&=\langle 0,0\rangle\end{cases}$ (20)
When $h=0$, the effective coupling between $\Phi$ and random strain vanishes, leading to no renormalization of $\gamma$, so $\gamma(h)=\Gamma$. When $h=1$, the effective random field width $w(h)$ reaches its maximum value of $W$. Finally, as $h\to\infty$ the enhanced interaction suppresses $\gamma$ and $w$ and the order parameter saturates to unity in the fully ordered state corresponding to $w=\gamma=t=0$.
Appropriate for the case of infinite-range interactions, we solve the RF-TFIM of Eq. (12) at $T\geq 0$, following the mean-field approach of [113]. We set $\tau_{i}^{z}=\Phi+\delta\Phi_{i}$ where $\Phi=\langle\tau_{i}^{z}\rangle$ and $\delta\Phi_{i}=\tau_{i}^{z}-\langle\tau_{i}^{z}\rangle$, and neglect the fluctuations, $\delta\Phi_{i}^{2}\approx 0$. This gives the mean field Hamiltonian
$\mathcal{H}_{\text{MF}}/\tilde{J}=\frac{N}{2}\Phi^{2}-\sum_{i}[(\Phi+z_{i})\tau_{i}^{z}+\gamma\tau_{i}^{x}]$ (21)
We compute the partition function $\mathcal{Z}[z_{i}]$ by employing the identity $\operatorname{Tr}e^{a\tau^{z}+b\tau^{x}}=2\cosh\sqrt{a^{2}+b^{2}}$ for real numbers $a$ and $b$, and we assume each disorder realization $\{z_{i}\}$ to consist of independent and identically distributed Gaussian random variables drawn from the distribution
$p(z_{i})\equiv\frac{1}{\sqrt{2\pi}w}e^{-z_{i}^{2}/2w^{2}}$ (22)
For each disorder realization, we compute the dimensionless free energy density $f=-\frac{T}{\tilde{J}N}\log\mathcal{Z}[z_{i}]$
$f(\Phi)=\frac{1}{2}\Phi^{2}-\frac{t}{N}\sum_{i}\log\Bigg{[}2\cosh\frac{\sqrt{(\Phi+z_{i})^{2}+\gamma^{2}}}{t}\Bigg{]}$ (23)
Assuming replica symmetry, we subsequently disorder-average $f\to\bar{f}$ over the joint distribution $\prod_{i}p(z_{i})$ and obtain the Landau expansion
$\bar{f}(\Phi)$ $=\frac{1}{2}\Phi^{2}-\int_{-\infty}^{\infty}p(z-\Phi)\Lambda(z)\mathrm{d}z$ (24)
$=\bar{f}(0)+\frac{\mathcal{A}}{2}\Phi^{2}+\frac{\mathcal{U}}{4}\Phi^{4}+\frac{\mathcal{G}}{6}\Phi^{6}+...$ (25)
In Eq. (24), we have defined the non-negative function $\Lambda(z)$ which encapsulates all of the $t(h)$ and $\gamma(h)$ dependence of the problem
$\Lambda(z)=t\log\Bigg{[}2\cosh\frac{\sqrt{z^{2}+\gamma^{2}}}{t}\Bigg{]}$ (26)
By expanding the integrand in powers of $\Phi$, the quadratic and quartic Landau coefficients can be written compactly as
$\mathcal{A}$ $=1-\int_{-\infty}^{\infty}\Lambda(z)\partial_{z}^{2}p(z)\mathrm{d}z$ (27)
$\mathcal{U}$ $=-\frac{1}{6}\int_{-\infty}^{\infty}\Lambda(z)\partial_{z}^{4}p(z)\mathrm{d}z$ (28)
Since $p(z)$ is Gaussian and since $\Lambda(z)$ grows algebraically, total derivative terms vanish. Consequently, one can “trade” $z$ derivatives between $p(z)$ and $\Lambda(z)$ at will using integration-by-parts. In the limit of zero temperature, $\Lambda$ simplifies dramatically:
$\lim_{T\to 0}\Lambda(z)=\lim_{t\to 0}\Lambda(z)=\sqrt{z^{2}+\gamma^{2}}$ (29)
Analytical expressions for the quadratic and quartic coefficients can be expressed using special functions:
$\mathcal{A}$ $=1-\frac{U(1/2,0,2x)}{\sqrt{2}w}$ (30)
$\mathcal{U}$ $=\frac{xe^{x}}{3\sqrt{2\pi}w^{3}}[(1+4x)K_{1}(x)-(3+4x)K_{0}(x)]>0$ (31)
where
$x\equiv\frac{\gamma^{2}}{4w^{2}}=\frac{\Gamma^{2}}{4\lambda^{2}H_{z}^{2}\bar{\varepsilon}_{s}^{2}}$ (32)
and $U(a,b,z)$ is a confluent hypergeometric function of the second kind, whereas $K_{n}(x)$ is a modified Bessel function of order $n$. Since $\mathcal{U}$ is strictly positive whenever $\mathcal{A}\geq 0$, the transition between AM and PM phases is second-order. Therefore, vanishing of the quadratic Landau coefficient $\mathcal{A}=0$ gives the phase boundary.
In the case of the classical RFIM with $\Gamma=0$ the quadratic Landau coefficient takes the form
$\mathcal{A}=1-\frac{U(1/2,0,0)}{\sqrt{2}w}=1-\frac{\sqrt{2/\pi}}{w}$ (33)
By solving $\mathcal{A}=0$ for $w$, we find the well-known result [100] for the critical disorder strength
$w_{c}=W_{0}\equiv\sqrt{2/\pi}\approx 0.798$ (34)
When $W>W_{0}$, AM reentrance occurs via a paramagnetic phase bounded by two field values given by $w(h_{c}^{\pm})=W_{0}$:
$h_{c}^{\pm}=\frac{W_{0}}{W}\Big{[}1\pm\sqrt{1-(W/W_{0})^{2}}\Big{]}$ (35)
For the case of nonzero $\Gamma$, the equation for the phase boundary $\mathcal{A}=0$ yields a critical disorder strength that decreases monotonically with $\Gamma$. We solve this equation and obtain the function $w_{c}(\gamma,t)$ restricted to the $t=0$ plane, i.e., $w_{c}(\gamma,0)$. We give the asymptotic behaviors near the clean QCP and classical disorder-induced transitions, respectively:
$w_{c}(\gamma,0)\approx\begin{cases}\sqrt{2/3}\sqrt{1-\gamma}&\text{for }\gamma\to 1\\ W_{0}\left(1+\frac{\pi}{4}\gamma^{2}\log\gamma\right)&\text{for }\gamma\to 0\end{cases}$ (36)
For fixed $\Gamma$, we may also obtain the minimum $W$ required to achieve reentrance. This minimum $W$, which we call $W_{*}$, is given by the condition of tangency between the $h$-parametrized trajectory $\boldsymbol{\theta}(W,\Gamma;h)$ and $\mathcal{A}=0$:
$\begin{cases}\mathcal{A}(\boldsymbol{\theta}(h))=0\\ \partial_{h}\mathcal{A}(\boldsymbol{\theta}(h))=0\end{cases}$ (37)
Solving these two equations yields $W_{*}$ and the magnetic field scale $h_{*}$ at which tangency occurs as explicit functions of $\Gamma$. Note that $W_{*}(\Gamma)$ and $h_{*}(\Gamma)$ are related via $w_{c}$ as an identity through the functional equation
$\frac{2h_{*}(\Gamma)}{1+h_{*}^{2}(\Gamma)}W_{*}(\Gamma)=w_{c}\Big{(}\frac{\Gamma}{1+h_{*}^{2}(\Gamma)},0\Big{)}$ (38)
We plot $W_{*}(\Gamma)$ and $h_{*}(\Gamma)$ in Fig. 4a-b, showing that the threshold disorder strength $W_{*}(\Gamma)$ and threshold field $h_{*}(\Gamma)$ fall monotonically with increasing $\Gamma$. The values at the classical limit $\Gamma=0$ and at the QCP $\Gamma=1$ are, respectively
$\begin{cases}W_{*}(0)=W_{0}=\sqrt{2/\pi}\approx 0.798\\ h_{*}(0)=1\\ W_{*}(1)=1/\sqrt{6}\approx 0.408\\ h_{*}(1)=0\end{cases}$ (39)
Interestingly, at $\Gamma=1$, $W_{*}$ is not zero, and instead approaches a minimum value of $W_{*}^{\mathrm{min}}=1/\sqrt{6}\approx 0.408$. This implies that from the definition of $W$, reentrance requires the energy scale of the random strain $E_{\text{el}}=\frac{1}{2}C_{0}\bar{\varepsilon}_{s}^{2}$ to be of order $J$ even at the QCP, where quantum disordering effects are the strongest. Conversely, the magnetic fields at which reentrance occurs are suppressed rapidly close to the QCP because $h_{*}\to 0$ as $\Gamma\to 1$.
Figure 4: (a) Threshold disorder strength $W_{*}/W_{0}$ and (b) corresponding magnetic field value $h_{*}$ as functions of $\Gamma$ at $T=0$. $W_{*}/W_{0}$ decreases closer to the QCP $\Gamma=1$ but remains of order one, whereas $h_{*}$ vanishes rapidly. This means that AM reentrance and order parameter non-monotonicity as a function of $h$ are favored when starting close to the zero magnetic field critical point.
The key advantage of representing the zero temperature phase diagram in the scaled coordinates $\gamma$ and $w$, as in Fig. 3(a), is that it does not explicitly depend on $W$ and $\Gamma$ and only does so implicitly through $w(h)$ and $\gamma(h)$. The drawback is that different $W$ correspond to differently curved phase trajectories $\boldsymbol{\theta}$. Thus, this representation may not be the most obvious physically motivated one and could obscure global features of the AM-PM phase boundary without a direct quantitative calculation. In some cases, it is more revealing to construct the zero-temperature phase diagram in coordinates for which the experimentally tunable quantity $h$ appears as an independent axis. This ensures that tuning magnetic field amounts to following a straight line, rather than a curve, for a fixed value of $\Gamma$ and $W$, with the drawback being that the phase boundary in these new coordinates depends explicitly on $W$.
We plot these $(\Gamma,h)$ phase diagrams for increasing values of fixed $W$ in Fig. 3(b)-(e). In panels (b) and (c), $W<W_{0}$, implying that the phenomenon of reentrant AM does not occur in the classical regime of $\Gamma=0$. Nevertheless, there is an important distinction between these two cases: in panel (b), because $W<W_{*}^{\mathrm{min}}<W_{0}$, reentrant AM does not happen anywhere in the phase diagram – recall that $W_{*}^{\mathrm{min}}=1/\sqrt{6}\approx 0.512W_{0}$. In this case, all that the magnetic field can do is tune the system from the PM to the AM phase, but it cannot tune the AM transition to zero. In contrast, in panel (c), $W_{*}^{\mathrm{min}}<W<W_{0}$, and the PM-AM phase boundary shows a “bulge” signaling a reentrant AM phase. This bulge first develops at $(\Gamma_{*},h_{*})=(1,0)$ when $W=W_{*}^{\mathrm{min}}$ and moves rapidly upward and to the left as $W$ increases further toward $\Gamma_{*}\rightarrow 0$ and $h_{*}\rightarrow 1$. When $W=W_{0}=\sqrt{2/\pi}$, which is the case shown in panel (d), the tip of the bulge annihilates at the $\Gamma=0$ axis, resulting in a non-analytic behavior of the free energy, as we discuss in Section IV. Finally, when $W>W_{0}$, illustrated in panel (e), reentrance occurs across the entire phase diagram between the critical fields $h_{c}^{\pm}$ satisfying Eq. (35). The combination of these panels gives the schematic three-dimensional phase diagram of Fig. 2.
Figure 5: Mean-field TFIM phase diagram, corresponding to $W=0$ in our case, plotted against scaled axes $t\equiv T/(1+h^{2})$ and $\gamma\equiv\Gamma/(1+h^{2})$. The scaled critical temperature $t_{c}=\gamma/\tanh^{-1}\gamma$ decreases with increasing $\gamma$, vanishing at a quantum critical point at $\gamma=1$ (in green). Increasing $h\equiv H_{z}/H_{\lambda}$ amounts to following a straight-line trajectory from $(\gamma,t)=(\Gamma,T)$ to the origin $(\gamma,t)=(0,0)$. The system orders monotonically due to the coupling to elastic fluctuations but no coupling to random strain.
III.3 Finite temperature phase diagram
Figure 6: (a): Classical (i.e., $\Gamma=0$) phase diagram with scaled axes $t\equiv T/(1+h^{2})$ and $w\equiv 2Wh/(1+h^{2})$, where $T=0.8J$ sets the starting point of the phase trajectory along the $t$-axis and $W$ sets the maximum extent of the trajectory along the $w$-axis. Green and blue curves correspond to $W=0.9W_{0}$ and $W=1.05W_{0}$, respectively. Panels (b) and (c) show the phase diagrams with unscaled axes $h=H_{z}/H_{\lambda}$ and $T$ corresponding to $W=0.9W_{0}$ and $1.05W_{0}$, respectively. As the relative disorder $W$ is increased, the paramagnetic region grows via the formation of a bulge in the AM-PM transition line at $(T_{*},h_{*})$, which annihilates at $(0,1)$ when $W=W_{0}=\sqrt{2/\pi}\approx 0.798$.
We now construct the finite temperature phase diagram in the four dimensional parameter space spanned by ($h,W,\Gamma,T$). As with the $T=0$ case, we represent the phase diagram in two different coordinate systems. The first is a 3D phase diagram with $h$-dependent coordinates ($w,\gamma,t$), where a changing magnetic field amounts to following a trajectory $\boldsymbol{\Theta}(h)$ in this space:
$\boldsymbol{\Theta}(W,\Gamma,T;h)=\langle w(h),\gamma(h),t(h)\rangle=\frac{\langle 2hW,\Gamma,T\rangle}{1+h^{2}}$ (40)
Note the special values
$\begin{cases}\boldsymbol{\Theta}(0)&=\langle 0,\Gamma,T\rangle\\ \boldsymbol{\Theta}(1)&=\langle W,\Gamma/2,T/2\rangle\\ \boldsymbol{\Theta}(\infty)&=\langle 0,0,0\rangle\end{cases}$ (41)
As in the case of zero temperature, $W$ represents the maximum extent of this trajectory along the $w$-axis.
Before going further, we consider two limiting cases for $T\geq 0$: the pure transverse field Ising model TFIM with $W=0$ and the classical RFIM with $\Gamma=0$. In both cases, $T_{c}$ decreases monotonically with increasing $\Gamma$ and $W$, resulting in QCPs at $\Gamma=1$ and $W=W_{0}$, respectively. These two QCPs are connected by a line of QCPs (Fig. 3a) whose shape on the ($\Gamma,W$)-plane depends on $h$.
When treating the pure TFIM, $p(z_{i})$ in Eq. (22) becomes a delta function, and the equation $\mathcal{A}=0$ for the phase boundary simplifies to
$1=\Lambda^{\prime\prime}(0)=\frac{\tanh(\gamma/t)}{\gamma}=(1+h^{2})\frac{\tanh(\Gamma/T)}{\Gamma}$ (42)
This yields the equation on $T_{c}$:
$T_{c}(\Gamma,h)=\frac{\Gamma}{\tanh^{-1}\frac{\Gamma}{1+h^{2}}}$ (43)
or, equivalently, on $t_{c}=\frac{T_{c}}{1+h^{2}}$:
$t_{c}(\gamma)=\frac{\gamma}{\tanh^{-1}\gamma}$ (44)
In this case without disorder, $\Phi$ couples only to elastic fluctuations, and increasing the magnetic field can only further order the system. This agrees with Eq. (43), from which we see that increasing $h$ suppresses the denominator and enhances $T_{c}$. The TFIM phase diagram represented in the ($t,\gamma$)-plane gives a convex curve – shown in Fig. 5 – with the $h$-parameterized trajectory $\boldsymbol{\Theta}(h)$ being a straight line connecting the initial point to the origin:
$\boldsymbol{\Theta}(0,\Gamma,T;h)=\frac{\langle\Gamma,T\rangle}{1+h^{2}}$ (45)
Therefore, in the clean case, the magnetic field can be used to tune the system from the PM phase to the AM phase, but not the other way around.
We now consider the classical RFIM case, for which $\Gamma=0$. In this limit, $\Lambda(z)$ simplifies, such that $\mathcal{A}=0$ gives
$t=\int_{-\infty}^{\infty}\,p(z)\operatorname{sech}^{2}(z/t)\mathrm{d}z$ (46)
Solving for $w$ yields the function $w_{c}(\gamma,t)$ along the $\gamma=0$ plane which has asymptotic behaviors given by
$w_{c}(0,t)\approx\begin{cases}\sqrt{1-t}&\text{for }t\to 1\\ W_{0}\big{(}1-\frac{\pi^{2}}{24}t^{2}\big{)}&\text{for }t\to 0\end{cases}$ (47)
Temperature and transverse field take qualitatively the same role. Both suppress $\Phi$ and both are renormalized in the same way due to elastic fluctuations. Therefore, we can obtain $W_{*}(T)$ analogously to how we obtained $W_{*}(\Gamma)$ in the previous section by determining the point of tangency on the ($w,t$)-plane, as shown in Fig. 6(a):
$\begin{cases}\mathcal{A}(\boldsymbol{\Theta}_{\Gamma=0}(h))=0\\ \partial_{h}\mathcal{A}(\boldsymbol{\Theta}_{\Gamma=0}(h))=0\end{cases}$ (48)
Solving these two equations yields $W_{*}$ and the magnetic field scale at which tangency occurs, $h_{*}$, as explicit functions of $T$. Note that $W_{*}(T)$ and $h_{*}(T)$ are related via $w_{c}$ through the functional equation
$\frac{2h_{*}(T)}{1+h_{*}^{2}(T)}W_{*}(T)=w_{c}\Big{(}0,\frac{T}{1+h_{*}^{2}(T)}\Big{)}$ (49)
As expected, the threshold scale falls monotonically with increasing $T$:
$\begin{cases}W_{*}(0)=W_{0}=\sqrt{2/\pi}\approx 0.798\\ h_{*}(0)=1\\ W_{*}(1)=1/2\\ h_{*}(1)=0\end{cases}$ (50)
Fig. 6(b)-(c) show two phase diagrams in unscaled coordinates $h$ and $T$ for representative values of disorder strength $W$.
The so far three convex phase boundaries that we have uncovered in the ($w,\gamma$)-, ($\gamma,t$)-, and ($w,t$)- planes suggest a convex critical surface in the ($w,\gamma,t$) parameter space. As with the 2D phase diagrams represented with $h$-dependent axes, the phase trajectory $\boldsymbol{\Theta}$ may avoid, intersect, or run tangent to this convex surface depending on the starting parameters $W$, $\Gamma$, and $T$.
Similarly to the $T=0$ and $\Gamma=0$ cases, we apply the condition for tangency on the whole ($w,\gamma,t$) parameter space, namely,
$\begin{cases}\mathcal{A}(\boldsymbol{\Theta}(h))=0\\ \partial_{h}\mathcal{A}(\boldsymbol{\Theta}(h))=0\end{cases}$ (51)
This yields the threshold values $W_{*}$ and $h_{*}$ as functions of $\Gamma$ and $T$, which, at finite temperatures, have a qualitatively similar behavior than the $T=0$ case shown in Fig. 4. We note that $W_{*}(\Gamma,T)$ remains of order one and is never suppressed to zero, suggesting that $E_{\text{el}}$ must be of order $J$ for reentrance to occur, like in the zero temperature case. However, we stress that $h_{*}$ can be made arbitrarily small depending on how close ($\Gamma,T$) is to the transition when $H_{z}=0$. Generalizing from the two-dimensional cases, $h_{*}(\Gamma,T)$ and $W_{*}(\Gamma,T)$ are related through the functional equation
$\frac{2h_{*}}{1+h_{*}^{2}}W_{*}=w_{c}\Big{(}\frac{\Gamma}{1+h_{*}^{2}},\frac{T}{1+h_{*}^{2}}\Big{)}$ (52)
Figure 7: Mean-field RF-TFIM phase diagram of an inhomogeneous $d$-wave altermagnet in ($w,\gamma,t$)-space. Shown as gray directed curves are three $H_{z}$-parameterized trajectories $\bf{\Theta}$ starting at fixed ($W,\Gamma$)=($J,0.75J$) for $T=0$, $0.5J$, and $J$.
Figure 8: Mean-field phase diagram of an inhomogeneous $d$-wave altermagnet as a function of magnetic field $H_{z}$, transverse field $\Gamma$ (responsible for quantum fluctuations), and temperature $T$ for different values of relative disorder strength $W=0.6J$ (a), $W=0.7J$ (b), and $W=0.8J>W_{0}$ (c). For $W>W_{0}$, the single AM phase region splits into two at $(H_{z},\Gamma,T)=(H_{\lambda},0,0)$ (dashed line).
In Fig. 7, we plot the critical surface in these scaled coordinates and show several phase trajectories. As with the previous section, we also display the critical surface in the ($h,\Gamma,T$) parameter space for different values of $W$. In Fig. 8, we see that when
$W>\max_{\Gamma,T}W_{*}(\Gamma,T)=W_{0}$ (53)
the single altermagnetic phase region separates at $(h,\Gamma,T)=(1,0,0)$ into two disconnected AM phase regions, one at low fields $h<1$ and one at high fields $h>1$. In this $W$ regime, altermagnetic reentrance is guaranteed to occur even at $T=0$.
IV Experimental signatures
The phase diagrams shown in Fig. 8 reveal that, in an inhomogeneous altermagnet, application of a magnetic field can induce both a PM-AM transition, when the disorder strength is weak, as well as an AM-PM-AM transition with reentrant altermagnetic order, when the disorder strength is strong. In this section, we go beyond the determination of the phase transition boundaries and discuss the behavior of three experimentally observable quantities as the phase boundaries are traversed by a magnetic field: the altermagnetic order parameter, the shear modulus, and the elasto-caloric effect coefficient. Note that, in principle, the AM order parameter could be measured in momentum space from the spin splitting of the band structure, or from extracting response functions that are proportional to the order parameter, such as the piezomagnetic and the elasto-Hall conductivity tensors [46].
IV.1 Altermagnetic order parameter
We first consider ($\Gamma,T$) values that place the system inside the AM ordered phase in the absence of a magnetic field. We also specialize to the case of $W<w_{c}(\Gamma,T)$ for which long-range AM order remains intact for all values of $h$. Although an AM reentrant behavior does not take place, we show that $\Phi(h)$ displays a non-monotonic behavior and a minimum value $\Phi_{\text{min}}$ for $h=h_{\text{min}}$.
The mean field equation $\partial\bar{f}/\partial\Phi=0$ gives a self-consistent condition for the order parameter:
$\Phi$ $=\int_{-\infty}^{\infty}p(z-\Phi)\partial_{z}\Lambda(z)\mathrm{d}z$ (54)
where we recall that $p(z)$ is a normal distribution with width $w=w(h)$ and
$\partial_{z}\Lambda(z)=\frac{z}{\sqrt{z^{2}+\gamma^{2}}}\tanh\frac{\sqrt{z^{2}+\gamma^{2}}}{t}$ (55)
In the limit of $h=0$, for which $(w,\gamma,t)=(0,\Gamma,T)$, a nonzero $\Phi$ satisfies the equation
$\sqrt{\Phi^{2}+\Gamma^{2}}=\tanh\frac{\sqrt{\Phi^{2}+\Gamma^{2}}}{T}$ (56)
for which we establish that the argument $\sqrt{\Phi^{2}+\Gamma^{2}}$ is less than unity due to the concavity of the hyperbolic tangent function. Conversely, the limit $h\to\infty$ yields a maximally ordered state for which $(w,\gamma,t)=(0,0,0)$. In this case, the self-consistent equation simply yields $|\Phi|=1$, i.e., the order parameter saturates to its maximum value. Finally, because $w(h)\propto h$ and $t(h),\gamma(h)\approx\text{const}$ for small $h$, increasing $h$ amounts to merely increasing the effective random field width $w$, which can only suppress $\Phi$. These three facts together, namely that $\Phi(h=0)\leq 1$, that $\Phi(h)$ decreases for small $h$, and that $\Phi(h=\infty)=1$, establishes that $\Phi$ must have a minimum at some magnetic field value $h=h_{\text{min}}$. It is straightforward to conclude that $h_{\text{min}}$ must be less than $1$, because an upturn in $\Phi$ can only happen when a nontrivial competition between the ordering and disordering tendencies occurs, i.e., when the effective energy scales $w(h)$, $\gamma(h)$, and $t(h)$ are not all decreasing, which can only happen if $h<1$. Generally, we can then write
$\Phi=\Phi_{\text{min}}+\frac{1}{2}\Upsilon(h-h_{\text{min}})^{2}+...$ (57)
where the combination of the self-consistent equation and the condition $d\Phi/dh=0$ yields a system of equations that give $h_{\text{min}}=h_{\text{min}}(W,\Gamma,T)$ and $\Phi_{\text{min}}=\Phi_{\text{min}}(W,\Gamma,T)$. We stress, however, that in the case of $\Phi=\mathcal{O}(1)$ deep inside the ordered phase, the necessity of minimizing the infinite-order expression for $\bar{f}(\Phi)$ yields a $\Phi_{\text{min}}=\Phi(h=h_{\text{min}})$ that is exponentially close to $\Phi(h=0)$, i.e.,
$|\Phi(h=0)-\Phi(h=h_{\text{min}})|\sim e^{-1/W^{2}h_{\text{min}}^{2}}$ (58)
In light of this fact, it is convenient to focus on the case where the parameter values place the system close to the AM-PM phase boundary, where $\bar{f}$ may be truncated to quartic order in $\Phi$. To proceed, we consider phase diagrams like those of Fig. 8(a)-(b), corresponding to $W<W_{0}$ (weak disorder). In these cases, reentrant behavior as a function of $h$ can be obtained by tuning the value of $\Gamma$. For concreteness, we focus on the $T=0$ cut of the phase diagram in the vicinity of the point $(\Gamma_{*},h_{*})$ that marks the end of the bulge of the PM-AM phase boundary, as shown in the inset of Fig. 9.
To obtain the form of $\Phi(h)$ as the PM-AM phase boundary is traversed upon increasing $\Gamma$, we use the phenomenological Landau coefficients. Since $\Phi$ is small, we can approximate $\mathcal{U}$ to be a constant and equal to $\mathcal{U}_{0}=\mathcal{U}(h=h_{\text{min}})$. Then, because $\mathcal{A}$ is analytic in $h$ and closest to zero when $h=h_{\text{min}}$, we have the approximation
$\mathcal{A}=-\mathcal{A}_{0}-\mathcal{A}_{2}(h-h_{\text{min}})^{2}$ (59)
where $\mathcal{A}_{0},\mathcal{A}_{2}>0$ are assumed to be $h$-independent constants. This yields
$\Phi=\sqrt{\frac{|\mathcal{A}|}{\mathcal{U}}}\approx\sqrt{\frac{\mathcal{A}_{0}+\mathcal{A}_{2}(h-h_{\text{min}})^{2}}{\mathcal{U}_{0}}}$ (60)
which gives a crossover field scale
$h_{t}\equiv\sqrt{\frac{\mathcal{A}_{0}}{\mathcal{A}_{2}}}$ (61)
that distinguishes between a quadratic and a linear dependence of the AM order parameter with respect to the distance to $h_{\mathrm{min}}$:
$\Phi(h)\propto\begin{cases}h_{t}\Big{[}1+\frac{(h-h_{\text{min}})^{2}}{2h_{t}^{2}}\Big{]}&\text{for }|h-h_{\text{min}}|\ll h_{t}\\ |h-h_{\text{min}}|&\text{for }h_{t}\ll|h-h_{\text{min}}|\ll 1\end{cases}$ (62)
If the tuning parameter $\Gamma$ is increased, we eventually encounter the case in which the PM-AM phase boundary is touched tangentially. In this case, $\mathcal{A}_{0}\rightarrow 0$ implying that $\Phi_{\min}\to 0$ and that the crossover scale $h_{t}$ vanishes, leaving a singular cusp-like behavior in the $h$-dependence of the order parameter for which long-range order is lost at precisely one value of magnetic field $h_{*}$:
$\Phi\propto|h-h_{*}|$ (63)
Note that, in the last step, we renamed $h_{\text{min}}$ to $h_{*}$ in accordance with the notation of Sec. III. These two behaviors for $\Phi(h)$, corresponding to a quadratic minimum and to a cusp are illustrated by the green and blue lines in Fig. 9 .
Figure 9: The AM order parameter $\Phi$ plotted as a function of the magnetic field $h$ at $T=0$ and $W=0.9W_{0}$ for $\Gamma=\Gamma_{*}(1-10^{-2})$ (green), $\Gamma=\Gamma_{*}$ (blue), and $\Gamma=\Gamma_{*}(1+10^{-2})$ (red), where $\Gamma_{*}\approx 0.526$. The associated phase diagram (inset) shows the corresponding line cuts in the magnified $T=0$ plane. For $\Gamma<\Gamma_{*}$, the order parameter displays an analytic local minimum at $h=h_{\text{min}}$, crossing over between quadratic (near $h_{\text{min}}$) and linear (away from $h_{\text{min}}$) regimes. For $\Gamma=\Gamma_{*}$, $h_{\text{min}}$ approaches $h_{*}\approx 0.842$ at which point $\Phi$ vanishes, and the quadratic regime is suppressed completely. For $\Gamma>\Gamma_{*}$, $\Phi$ is singular at the two critical fields $h_{c}^{\pm}$, crossing over between square-root (near $h_{c}^{\pm}$) and linear (away from $h_{c}^{\pm}$) regimes.
Figure 10: The zero-temperature mean-field order parameter as a function of $h=H_{z}/H_{\lambda}$ for different values of $\Gamma$ given in the legend, and $W=0.9W_{0}$ (d), $W=W_{0}$ (e), and $W=1.05W_{0}$ (f). The corresponding cuts in the $(h,\Gamma)$-plane are given in panels (a), (b), and (c), respectively. For weak disorder, the order parameter behaves non-monotonically, whereas for strong disorder it exhibits reentrance.
Upon increasing $\Gamma$ further, the PM-AM phase boundary is intersected twice at the two critical fields $h_{c}^{\pm}$, signaling the loss of long-range order over a finite range $h\in[h_{c}^{-},h_{c}^{+}]$. In the limit of being near the intersection point, i.e., for $h\to h_{c}^{-}$ or $h\to h_{c}^{+}$, the phase trajectory $\boldsymbol{\Theta}(h)$ has a nonzero component normal to the phase boundary, and hence the distance to the critical surface scales linearly with $|h-h_{c}^{\pm}|$. For $\Phi(h)$, this translates to a “splitting” of the cusp at $h_{*}$ into two successive square-root singularities at $h_{c}^{-}$ and $h_{c}^{+}$.
We use our Landau theory to investigate what happens within the ordered phase when $h_{c}^{-}$ and $h_{c}^{+}$ are sufficiently close together. Within the ordered phase, but near $h_{c}^{\pm}$, we obtain for the quadratic coefficient
$\mathcal{A}=-\mathcal{A}_{0}^{\prime}(h-h_{c}^{-})(h-h_{c}^{+})$ (64)
where $\mathcal{A}_{0}^{\prime}>0$ such that $h\in[h_{c}^{-},h_{c}^{+}]$ yields $\mathcal{A}>0$ and $h\notin[h_{c}^{-},h_{c}^{+}]$ yields $\mathcal{A}<0$. Using $\Phi\sim\sqrt{-\mathcal{A}/\mathcal{U}}$ we obtain a crossover regime governed by
$h_{t}^{\prime}\equiv h_{c}^{+}-h_{c}^{-}$ (65)
that distinguishes between square-root and linear behaviors:
$\Phi(h)\propto\begin{cases}\sqrt{h_{t}^{\prime}}\sqrt{h_{c}^{-}-h}&\text{for }0<h_{c}^{-}-h\ll h_{t}^{\prime}\\ \sqrt{h_{t}^{\prime}}\sqrt{h-h_{c}^{+}}&\text{for }0<h-h_{c}^{+}\ll h_{t}^{\prime}\\ |h-(h_{c}^{+}+h_{c}^{-})/2|&\text{for }h_{t}^{\prime}\ll|h-h_{c}^{\pm}|\ll 1\end{cases}$ (66)
This behavior is illustrated by the red line in Fig. 9. We can also decrease $\Gamma$ to move back from the reentrant to the tangent case. The corresponding limit is $h_{c}^{\pm}\to h_{*}$ and $h_{t}^{\prime}\to 0$, for which the square-root behavior vanishes and only the cusp-like behavior of Eq. (63) remains. We note that Eq. (66) is an approximation, and higher-order corrections in $h_{t}^{\prime}$ will lead to square-root prefactors that are different for the two critical fields.
While the analysis here focused on the $T=0$ phase diagram with weak disorder ($W<W_{0}$), similar behaviors for $\Phi(h)$ are carried over to the phase diagrams of the marginal disorder ($W=W_{0}$) and strong disorder ($W>W_{0}$) cases. As shown in Fig. 10, in the former, $\Phi(h)$ displays the cusp-behavior for $\Gamma=0$ whereas, in the latter, $\Phi(h)$ always shows the square-root-like behavior for any $\Gamma$ value.
IV.2 Shear modulus softening
While in the previous section we focused on the behavior of the AM order parameter, in this and in the next subsection we focus on the behavior of the AM susceptibility. The key point is that piezomagnetism enables a coupling between altermagnetic and elastic degrees of freedom such that their fluctuations also become correlated. Consequently, order parameter fluctuations, which become pronounced near the AM transition, enhance elastic fluctuations and lead to a structural transition which accompanies the AM critical point (see also [114]). The situation is analogous to a nematic transition softening the shear modulus [115]. A straightforward mean field calculation of the renormalized shear modulus $\tilde{C}_{0}$ via Eq. (6) gives:
$\tilde{C}_{0}^{-1}=C_{0}^{-1}+\frac{\lambda^{2}H_{z}^{2}}{C_{0}^{2}}\chi$ (67)
where $\chi$ is the AM static susceptibility. As expected, the $H_{z}=0$ limit results in no renormalization since the effective bilinear strain-AM coupling is zero, whereas $H_{z}\neq 0$ results in the vanishing of the shear modulus at the AM transition where $\chi\to\infty$. Using the definition of $H_{\lambda}$, Eq. (67) can be written in a dimensionless fashion:
$\frac{\tilde{C}_{0}}{C_{0}}=\frac{1}{1+h^{2}\chi}$ (68)
This analysis reveals a clear analogy between nematicity and altermagnetism [114]. In nematicity, a direct bilinear coupling between the order parameter and shear strain is symmetry-allowed, resulting in a structural distortion at the nematic transition. This coupling of an electronic order parameter to the lattice, and subsequent structural distortion, is a key ingredient in the elucidation of nematicity in correlated systems [116, 117]. By contrast, in the altermagnetic case, this effective coupling and the resulting structural distortion magnitude is tunable by a magnetic field.
To calculate the renormalized shear modulus, we use our Landau theory to ascertain $\chi$ in both the PM and AM phases. In mean field theory, $\chi$ exhibits Curie-Weiss-like universal behavior proximate to the phase boundary in which
$\chi\propto|x-x_{c}|^{-1}$ (69)
Figure 11: The altermagnetic susceptibility $\chi\equiv[\partial\Phi/\partial b]_{b\to 0}$ (a) and its inverse (b) plotted as a function of $h$ at $T=0$ and $W=0.9W_{0}$ for different values of $\Gamma$, as shown in the inset phase diagram. Here, $\Gamma_{*}\approx 0.526J$ and $h_{*}\approx 0.842$. The green curve is the susceptibility for the non-reentrant case $\Gamma=\Gamma_{*}(1-10^{-2})$, for which $\chi$ achieves a quadratic maximum at $h=h_{\text{min}}$. In blue, we plot $\chi$ in the tangent case $\Gamma=\Gamma_{*}$ for which $\chi$ diverges as $|h-h_{*}|^{-2}$. In red, we plot $\chi$ in the reentrant case $\Gamma=\Gamma_{*}(1+10^{-2})$, for which $\chi$ diverges as $|h-h_{c}^{\pm}|^{-1}$ at the two successive critical fields.
Here, $x$ is a generic tuning parameter parameterizing an axis with a component that is normal to the phase boundary marked by $x_{c}$. As with $\Phi$, the various trajectories followed on the phase diagram with different choices of $\Gamma$, $W$, and $T$ leads to interesting consequences on the behavior of $\chi$ as a function of $h$, such as non-monotonicity or a suppression of the universal regime arbitrarily close to the critical field. To calculate the AM susceptibility, we introduce an infinitesimal conjugate field $b$ which couples to $\sum_{i}\tau_{i}^{z}$ in the Hamiltonian. To take into account the fact that we have divided the Hamiltonian through by the renormalized exchange $\tilde{J}=J[1+h^{2}]$, we define
$B(h)\equiv\frac{b}{1+h^{2}}$ (70)
In the Landau theory, $B$ simply shifts the mean of the effective random field distribution $p(z)$. This yields the new free energy
[…]
To lowest order in $B$, the saddle-point equation $\partial_{\Phi}\bar{f}_{B}=0$ can be written as
$\partial_{\Phi}\bar{f}=B(1-\partial_{\Phi}^{2}\bar{f})$ (72)
where only the $B=0$ free energy $\bar{f}$, defined in Eq. (24), is invoked. Differentiating Eq. (72) with respect to $b$ and taking $b\to 0$ yields the susceptibility in terms of $\bar{f}$ or, equivalently, its Landau coefficients $\mathcal{A}$ and $\mathcal{U}$:
$\chi$ $=\frac{[\partial_{\Phi}^{2}\bar{f}]^{-1}-1}{1+h^{2}}$ (73)
$=\frac{1}{1+h^{2}}\times\begin{cases}\mathcal{A}^{-1}-1&\text{for}\Phi=0\\ {}[\mathcal{A}+3\mathcal{U}\Phi^{2}+...]^{-1}-1&\text{for }\Phi\neq 0\end{cases}$ (74)
Using this expression for $\chi$, the shear modulus renormalization can be written as
$\frac{\tilde{C}_{0}}{C_{0}}=\frac{1+h^{2}}{\partial_{\Phi}^{2}\bar{f}+h^{2}}(\partial_{\Phi}^{2}\bar{f})$ (75)
In Fig. 11, we plot the behavior of the field-dependent AM susceptibility $\chi(h)$ and of its inverse along different cuts in the zero-temperature $(\Gamma,h)$ phase diagram shown in the inset. We focus on the regime of weak disorder $W_{*}<W<W_{0}$, for which reentrant AM order emerges to the right of the $(\Gamma_{*},h_{*})$ point that marks the end of the bulge of the PM-AM phase boundary. The non-monotonic behavior of $\chi$ is evident, as is the unusual quadratic behavior of $\chi^{-1}$ when the PM-AM phase boundary is tangentially touched (blue curve).
The temperature dependence of the renormalized shear modulus $\tilde{C}_{0}/C_{0}$ is shown in Fig. 12 for different values of the magnetic field. Here, we consider the $(T,h)$ phase diagram cross section show in the inset, with fixed $W=0.7J$ and $\Gamma=0.3J$. Without a magnetic field, as expected, there is no change in the shear modulus. However, as the field increases, the shear modulus vanishes as the transition is approached, signaling a tetragonal-to-orthorhombic transition coincident with the PM-AM transition.
[…]
IV.3 The elasto-caloric effect
Another experimental quantity that depends on the AM susceptibility is the elasto-caloric effect (ECE) coefficient. The ECE can be thought of heat generation in response to an applied ac strain [118]. The precise measurement of this effect, namely the amount of heat arising from a given amount of strain, provides a direct probe of the entropy landscape and is therefore sensitive to phase transitions. Indeed, the ECE has been exploited to detect strain-tunable phase transitions in nematic [119, 120], superconducting [121, 122, 123], and magnetic systems [124, 125]. Due to the nontrivial role of strain in AM systems, ECE provides a powerful tool for the experimental detection and characterization of AM phases and phase transitions.
The ECE coefficient $\eta$ – having units of energy – is defined as the strain derivative of temperature taken at constant entropy. Recall that in our units, all energy scales including $\eta$ are in units of the bare interaction $J$. The ECE coefficient is
$\eta=\Big{(}\frac{\partial T}{\partial\varepsilon}\Big{)}_{S}=-\frac{T}{C_{\varepsilon}}\Big{(}\frac{\partial S}{\partial\varepsilon}\Big{)}_{T}$ (76)
Here, $\varepsilon$ is a small applied shear strain. Note that $\eta$ can be thought of as a version of the Grüneisen parameter which connects $T$ to $\varepsilon$. The Maxwell relation $\Big{(}\frac{\partial S}{\partial\varepsilon}\Big{)}_{T}=-\Big{(}\frac{\partial S}{\partial T}\Big{)}_{\varepsilon}\Big{(}\frac{\partial T}{\partial\varepsilon}\Big{)}_{S}$ yields the second equality, where $C_{\varepsilon}=T\Big{(}\frac{\partial S}{\partial T}\Big{)}_{\varepsilon}$ is the heat-capacity at constant $\varepsilon$. We can equivalently work with specific entropy $s\equiv S/N$ and specific heat $c\equiv C_{\varepsilon}/N$ since a scaling of either does not change the result for $\eta$. Using our Landau theory, it is straightforward to compute the specific entropy $s=-\frac{\partial\bar{f}}{\partial T}$ via a partial derivative with respect to temperature rather than a total one, since we do not need to consider the temperature dependence in the order parameter as $\partial\bar{f}/\partial\Phi=0$.
To set the stage, and in order to properly connect the ECE coefficient to other physical observables like AM susceptibility and the shear modulus, we take account of the quantities held fixed in the definition of each observable. Recall that the order parameter susceptibility and the shear modulus are defined at fixed zero stress. This allowed us to treat the elastic degrees of freedom, i.e., $\varepsilon_{0}$ in Eq. (6), dynamically, thereby renormalizing $J\to\tilde{J}$. At nonzero $H_{z}$, dynamical elastic modes contributed to the ordering tendency that opposed the disordering tendency caused by random strain, resulting in the nontrivial phase diagrams found in previous sections.
On the other hand, the ECE is measured at fixed strain, with the ac strain period being longer than any inherent relaxation timescale. We therefore cannot treat elastic modes dynamically in this scenario and, in our model, no renormalization of $J$ occurs. As a result, any divergence in the expression for $\eta$ at nonzero $H_{z}$ occurs at a temperature $T^{*}$ which is strictly lower than $T_{c}$. In other words, the ECE coefficient depends on the bare AM susceptibility rather than the actual AM susceptibility. Consequently, $\eta$ does not diverge at the true AM transition and is only enhanced to a finite value, as we will show in the ECE coefficient curves of Fig. 13. Note that, while only the curves associated with $H_{z}=0.5H_{\lambda}$ and $H_{z}=0.3H_{\lambda}$ in panel (a) seem to terminate at a finite value, a larger vertical axis window would show that the other curves do as well.
We compute $\eta$ in the disordered phase, and expand our Landau free energy to quartic order to obtain its behavior near the transition $T_{c}$ (assuming that $h\lesssim 1$ so that $T^{*}$ and $T_{c}$ are not too far apart). An infinite-order analysis is done in the Appendix to obtain $\eta$ for all parameter values. As in the previous section, we introduce a conjugate field $b$, which we now assume to be arising from an applied infinitesimal shear strain $b=\lambda H_{z}\varepsilon$ and expand the free energy density to $\mathcal{O}(\Phi^{4})$:
$\bar{f}(\Phi)=\frac{\mathcal{A}}{2}\Phi^{2}+\frac{\mathcal{U}}{4}\Phi^{4}-b\Phi$ (77)
Crucially, the Landau parameters are given in Eq. (27) and Eq. (28) but with the important condition that $\tilde{J}=J$ due to the frozen out elastic modes. We can re-express $\mathcal{A}$ in terms of the temperature scale $T^{*}$ as
$\mathcal{A}=\mathfrak{a}(T-T^{*})$ (78)
where $\mathfrak{a}$ and $T^{*}$ generically depend on all non-thermal parameters in the problem. Taking $\mathcal{A}=0$ in the case of no random strain disorder $W=0$ yields
$T^{*}(\Gamma,h)=\frac{\Gamma}{\tanh^{-1}\Gamma}$ (79)
Using the expression for $T_{c}$ in Eq. (43), we find to leading order in $h$
$T_{c}-T^{*}\approx\frac{\Gamma^{2}h^{2}}{(1-\Gamma^{2})(\tanh^{-1}\Gamma)^{2}}$ (80)
which in the classical limit of $\Gamma\to 0$ becomes $T_{c}-T^{*}\approx\lambda^{2}H_{z}^{2}/C_{0}$ and recovers the well-known case of coupled nematic and structural degrees of freedom [126] (where we interpret $\lambda H_{z}$ as the effective nemato-elastic coupling).
From here, the specific entropy as a function of $\Phi$ follows as $s=-\partial\bar{f}/\partial{T}=-\mathfrak{a}\Phi^{2}/2$ where $\Phi$ solves $\partial\bar{f}/\partial\Phi=0$, yielding the expression
$s=-\frac{b^{2}}{2\mathfrak{a}(T-T^{*})^{2}}=-\frac{\lambda^{2}H_{z}^{2}}{2\mathfrak{a}(T-T^{*})^{2}}\varepsilon^{2}$ (81)
Hence, the ECE coefficient $\eta$ follows from Eq. (76) as
$\eta$ $=\frac{T}{c_{+}}\frac{\lambda^{2}H_{z}^{2}}{\mathfrak{a}(T-T^{*})^{2}}\varepsilon$ (82)
$=\frac{T}{c_{+}}\frac{\lambda^{2}H_{z}^{2}}{\mathfrak{a}(T-T_{c}+\varsigma\lambda^{2}H_{z}^{2}/C_{0})^{2}}\varepsilon$ (83)
where $c_{\text{+}}$ refers to the specific heat immediately above $T^{*}$ and $\varsigma$ is an $H_{z}$-independent positive dimensionless $\mathcal{O}(1)$ parameter. In accordance with our mean field approach, the specific heat undergoes a finite jump $\Delta c=c_{+}-c_{-}$ at $T=T^{*}$. In the limit of $H_{z}\ll H_{\lambda}$, the greatest contribution to the variation in $\eta$ arises from the temperature denominator. Accordingly, the maximum enhancement of $\eta$ occurs close to the AM transition, with a maximum value $\eta_{\text{max}}\propto\lambda^{2}H_{z}^{2}(T_{c}-T^{*})^{-2}$ that diverges in the limit $H_{z}\to 0$. This suggests a crossover temperature at small fields for which $\eta\sim(T-T_{c})^{-2}$:
$h^{2}\ll\frac{T-T_{c}}{J}\ll 1\implies\eta\propto\frac{\lambda^{2}H_{z}^{2}}{(T-T_{c})^{2}}\varepsilon$ (84)
Note that the $(T-T_{c})^{-2}$ divergence is cutoff when $T-T_{c}\sim Jh^{2}$.
For $h\sim 1$, $T_{c}$ and $T^{*}$ are generally quite different, and the quartic truncation for $\bar{f}$ is insufficient. An infinite-order expression for $\eta$ is derived in the Appendix. Using the equations derived in the Appendix, we obtain $\eta/C_{0}\varepsilon$ as a function of $T$ for fixed $\Gamma,W$ and several values of $h$, as plotted in Fig. 13. We clearly see that, as $h$ increases, the enhancement of the ECE coefficient $\eta$ at the AM transition becomes weaker. The behavior of $1/\sqrt{\eta}$ is to be contrasted with that of the renormalized shear modulus $\tilde{C}_{0}$ in Fig. 12 for the same magnetic field values in the same $(T,h)$ phase diagram. Specifically, $\tilde{C}_{0}$ always vanishes at the transition, as it is proportional to the actual AM susceptibility, whereas $1/\sqrt{\eta}$ always reaches a non-zero value at the AM transition, since it is proportional to the bare AM susceptibility.
Figure 13: ECE coefficient $\eta/C_{0}\varepsilon$ (a) and its inverse square root $\sqrt{C_{0}\varepsilon/\eta}$ (b) for different field values $h=H_{z}/H_{\lambda}$ as a function of $T$ for $W=0.7J$, $\Gamma=0.3J$ (the same as in Fig. 12, as shown by the phase diagram in the inset). All curves terminate at the AM phase transition. The ECE is enhanced near the AM transition, but remains finite, and exhibits a $|T-T_{c}|^{-2}$ scaling regime at modest fields for temperatures satisfying $Jh^{2}\ll T-T_{c}\ll J$ [Eq.(84)]. The dashed line is the AM transition temperature without a field.
V Beyond Mean field
In realistic systems, true infinite-range interactions do not exist and one must explicitly consider the effect of finite wavelength elastic fluctuations, which mediate long-range but not infinite-range interactions between AM degrees of freedom. In this case, the effect of fluctuations above mean-field become relevant for both bulk-phase quantities and in the universal behavior near the transition. The additional complication of random fields makes the RF-TFIM extremely difficult to solve beyond mean-field. Indeed, the effect of rare regions promotes activated universal scaling of dynamic observables, causing super-exponential growth of timescales near the phase transition [108, 127, 128]. Additionally, the generally slow-dynamics close to the transition makes Monte Carlo methods also quite resource-intensive. However, for some dynamical processes on not too long time scales and, in particular, for a qualitative understanding of some of the equilibrium properties, the mean field analysis discussed here still gives a sensible qualitative behavior.
Since we do not consider dynamics in this work, we assert that the qualitative behavior of observables obtained from the mean field theory should carry over to the general case. Moreover, the fact that the phase diagram is a deformation of the RF-TFIM phase diagram must hold in general, and not just in mean field.
VI Discussion and Conclusions
Piezomagnetism is a generic phenomenon of altermagnetic systems that enables the tunability of the effective AM-strain coupling by external magnetic fields. This trilinear coupling leads to competition between two opposing phenomena: random strain and long-wavelength elastic fluctuations. The former arises from crystalline defects, and, when multiplied by a magnetic field, acts as an effective random longitudinal field conjugate to the AM order parameter $\Phi$. Applying strong magnetic fields continuously enhances this random field strength linearly in $H_{z}$ and thereby pushes the system across the critical disorder strength $w_{c}$ of an effective RF-TFIM. On the other hand, long-wavelength elastic fluctuations enable long-range interactions between AM degrees of freedom, driving the system mean field, with the effective long-range coupling strengthening as $H_{z}^{2}$ and thus favoring an ordered state. Thus, at weak magnetic fields, the random strain effect dominates, whereas at strong magnetic fields, the elastic fluctuations effect dominates, with a crossover field set by $H_{\lambda}$. At this field strength, the effective strain-AM coupling $\lambda H_{\lambda}$ equals the geometric mean of the AM energy scale $J$ and the bare shear modulus $C_{0}$.
By minimizing the mean-field free energy of the RF-TFIM, we uncovered a nontrivial phase diagram as a function of effective random strain strength $W$ (proportional to $\sqrt{E_{\text{el}}/J}$), on-site tunneling amplitude $\Gamma$, temperature $T$, and magnetic field $H_{z}$. This phase diagram is simply a $W$-dependent deformation of the original convex critical surface of the RF-TFIM, where the axes are given by magnetic-field dependent effective temperature $t$, effective random field strength $w$, and effective transverse field $\gamma$.
For fixed $\Gamma$, $T$, there is a threshold disorder strength value $W_{*}(\Gamma,T)$ that distinguishes three types of behaviors once the system starts in the AM phase. For $W>W_{*}$, an increasing magnetic field first drives the AM phase towards a PM phase at $H_{z,c}^{-}<H_{\lambda}$ and then back to the AM phase at $H_{z,c}^{+}>H_{\lambda}$, establishing a reentrant AM order. Exactly at $W=W_{*}$, the critical fields $H_{z,c}^{\pm}$ merge onto a single point $H_{z,*}$ and the corresponding critical point has critical exponents that are twice the values of the critical exponents at the separate points $H_{z,c}^{\pm}$. For $W<W_{*}$, the system remains in the AM phase. In all cases, however, if the system starts in the PM phase, a large enough field may drive a transition to the AM phase.
We found that each of these different regimes leave clear fingerprints on the AM order parameter and the AM susceptibility. For instance, even in the case $W<W_{*}$, where the AM phase cannot be completely suppressed by an external magnetic field, the AM order parameter and the AM susceptibility show a non-monotonic behavior. Importantly, we showed that the AM susceptibility $\chi$ is directly manifested in the renormalized shear modulus $\tilde{C}_{0}$, which vanishes as $1/\chi$, and in the elasto-caloric effect coefficient $\eta$ , which enhances strongly near the transition in the presence of a magnetic field, thus opening new avenues to probe AM fluctuations in the PM phase.
We emphasize that the tuning properties of the external magnetic field become much more pronounced if the system is closer to the AM-PM phase boundary, in which case the critical fields $H_{z,c}^{\pm}$ become small. Nevertheless, it is instructive to estimate the typical field magnitude $H_{\lambda}$ that sets the scale for many of the phenomena discussed here to occur. Minimizing the free energy $F(\Phi,\varepsilon)$ of piezomagnetically coupled order parameter and strain with respect to uniform strain $\varepsilon$ yields the linear response relation $\varepsilon=\lambda\Phi H_{z}/C_{0}$, which allows us to directly identify the piezomagnetic tensor element $\Lambda=\lambda\Phi/C_{0}$. Note that $\Lambda$ has dimensions of inverse magnetic field, since $\Phi$ is dimensionless, $C_{0}$ has dimensions of energy, and $\lambda$ has dimensions of energy times inverse field. Assuming $\Phi=\mathcal{O}(1)$ deep inside the AM phase, this enables us to relate the magnetic field scale $H_{\lambda}$ with other quantities:
$H_{\lambda}=\frac{\sqrt{C_{0}J}}{\lambda}\sim\frac{1}{\Lambda}\sqrt{\frac{J}{C_{0}}}$ (85)
Now, the energy scale associated with the shear modulus $C_{0}$ is typically of the order of $30$ eV, corresponding to an elastic constant of the order of $\mathcal{C}_{0}\sim 50$ GPa multiplied by a unit cell volume of the order of $v\sim 100$3, i.e., $C_{0}=\mathcal{C}_{0}v$. Conversely, the energy scale $J$ can be estimated from the typical AM transition temperature, giving $10$ meV, and thus $H_{\lambda}\sim 0.02/\Lambda$ As for the piezomagnetic coupling $\Lambda$, we consider rutiles like CoF2, FeF2, and MnF2, which are $d_{xy}$-wave altermagnets. The quantity reported in the literature for CoF2 [129, 130], $\bar{\Lambda}=6\times 10^{-3}\mu_{B}/\mathrm{GPa}$ , is actually the magnetization induced by applied stress, $m_{z}=\bar{\Lambda}\sigma_{xy}$. Since $m_{z}=\chi_{0}H_{z}$ and $\sigma_{xy}=\mathcal{C}_{0}\varepsilon$, we thus have $1/\Lambda=\bar{\Lambda}\mathcal{C}_{0}/\chi_{0}\sim 6T$. In the last step, we approximated the uniform magnetic susceptibility to a characteristic value for these rutiles [131, 132], $\chi_{0}\sim 0.05\mu_{B}/T$. Putting everything together, we find $H_{\lambda}\sim 0.1T$, which is an accessible value. We can also use these estimates to express the disorder strength in terms of the root mean square strain $\bar{\varepsilon}_{s}$, obtaining $W=\bar{\varepsilon}_{s}\sqrt{C_{0}/4J}\sim 30\bar{\varepsilon}_{s}$. This implies that $W\sim 1$, which ensures the existence of reentrant AM behavior, requires $\bar{\varepsilon}_{s}\sim 3\%$, which is a reasonable value.
acknowledgments
We thank T. Birol, A. V. Chubukov, I. Jang, W. J. Meese, L. Šmejkal, and C. Steward for fruitful discussions. A.R.C. and R.M.F. were supported by the Air Force Office of Scientific Research under Award No. FA9550-21-1-0423. J.S. was supported by the German Research Foundation TRR 288-422213477 ELASTO-Q-MAT, Project A07. R.M.F. also acknowledges a Mercator Fellowship from the German Research Foundation (DFG) through TRR 288, 422213477 Elasto-Q-Mat.
Appendix A Derivation of the elasto-caloric effect coefficient
To obtain the ECE coefficient, we freeze out elastic modes and set $\tilde{J}=J$ in the infinite-order expression for $\bar{f}$ in Eq. (24). We let $b=\lambda H_{z}\varepsilon$ be an infinitesimal bias arising from an infinitesimal applied strain $\varepsilon$, and compute the specific entropy as
$s=-\frac{\partial\bar{f}}{\partial T}$ $=\int_{-\infty}^{\infty}p(z-\Phi-b)\partial_{t}\Lambda(z)\mathrm{d}z$ (86)
$=\log 2-\int_{-\infty}^{\infty}p(z-\Phi-b)\Big{[}\frac{\sqrt{z^{2}+\gamma^{2}}}{t}\tanh\frac{\sqrt{z^{2}+\gamma^{2}}}{t}-\log\cosh\frac{\sqrt{z^{2}+\gamma^{2}}}{t}\Big{]}\mathrm{d}z$ (87)
where
$p(z)\equiv\frac{1}{\sqrt{2\pi}w}e^{-z^{2}/2w^{2}}$ (88)
The term in the square brackets of the integrand in Eq. (87) is a non-negative function bounded above by $\log 2$, i.e., $s\leq\log 2$. Note that because $\tilde{J}=J$, the definitions for the scaled variables $t$, $\gamma$, and $w$ are modified slightly, namely, $t\equiv T$, $\gamma\equiv\Gamma$, and $w=2hW$. Using this new definition, it follows that
$\frac{\partial s}{\partial\varepsilon}$ $=\lambda H_{z}\Big{(}1+\frac{\partial\Phi}{\partial b}\Big{)}\int_{-\infty}^{\infty}p(z-\Phi-b)\frac{z}{T^{2}}\operatorname{sech}^{2}\frac{\sqrt{z^{2}+\Gamma^{2}}}{T}\mathrm{d}z$ (89)
$\frac{\partial s}{\partial T}$ $=\int_{-\infty}^{\infty}p(z-\Phi-b)\Big{(}\frac{\partial\Phi}{\partial T}+\frac{\sqrt{z^{2}+\Gamma^{2}}}{T}\Big{)}\frac{\sqrt{z^{2}+\Gamma^{2}}}{T}\operatorname{sech}^{2}\frac{\sqrt{z^{2}+\Gamma^{2}}}{T}\mathrm{d}z$ (90)
where $\Phi$ is the solution of $\partial\bar{f}/\partial\Phi=0$:
$\Phi=\int_{-\infty}^{\infty}p(z-\Phi-b)\frac{z}{\sqrt{z^{2}+\Gamma^{2}}}\tanh\frac{\sqrt{z^{2}+\Gamma^{2}}}{T}\mathrm{d}z$ (91)
We solve Eq. (91) numerically in the parameter space and combine Eq. (89) and Eq. (90) to obtain the ECE coefficient $\eta$ shown in Fig 13.